OpenStudy (anonymous):

solve x^-2x-13=0 using quadratic equation from india

OpenStudy (anonymous):

from india?

OpenStudy (anonymous):

no solutions

OpenStudy (anonymous):

i think haha

OpenStudy (anonymous):

\[x^2-2x-13=0\] \[x^2-2x=13\] \[(x-1)^2=13+1=14\] \[x-1=\pm\sqrt{14}\] \[x=1\pm\sqrt{14}\] is easiest method

OpenStudy (tyteen4a03):

Quadratic Formula is \[x = \frac{ -b \pm \sqrt{b^2 - 4ac } }{ 2a }\] I noticed a possible type and assumes that the equation is \(x^2-2x-13=0\). Just substitute a = 1, b = -2 and c = -13 into this formula. @satellite73 Indeed it is, but some questions require you to be familiar with the Quadratic Equation.

OpenStudy (anonymous):

@satellite73 got it. i just realized i multiplied negative numbers and got a negative answer ??

OpenStudy (anonymous):

Wait is there a different way to solve it in india?

OpenStudy (anonymous):

because it's from india.

OpenStudy (anonymous):

actually "formula from india" only had positive square root. it wasn't until 8th century that al-kwarithimi (whence both the word "al-gorism" and also al-jabr aka "algebra") gave the complete solution

OpenStudy (anonymous):

lol if they are not using x equals negetive b plus or minus radical b sqaured minus four a c all over two a then they are not getting the correct roots... hahaha thats all i can say...

OpenStudy (mayankdevnani):

\[x ^{2}-2x-13=0\] \[x ^{2}-2x=13\] \[x ^{2}-2x+1=13+1\] \[x ^{2}-2x+1=14\] Factors Left- \[(x-1)\times(x-1)=14\] \[(x-1)^{2}=14\] \[x-1=\sqrt{14}\] \[x=\sqrt{14}+1\]